6. Given: Mass 0.106 kg, Mass2 0.121 kg 2. Acceleration 0.585 m/s g 9.8 m/s2 What is the calculated applied force? (Neglect friction.) 0.133 (answer with three sig figs) 7. Given: Mass 0.111 kg; Mass 0.116 kg; Acceleration 0.202 m/s g 9.8m/s 2 What is the theoretical Newtons Second Law Force? (Neglect friction.) (note: Use Three Significant Figures or N/A if there is no answer.)
Answer #1

according to Newton's second law

6)

F=F_{1}+F_{2}\\ \\ F=0.106\times 0.585+0.121\times 0.585\\ \\ F=0.133N

+++++++++++++++++++++++++++++++++++++++++

7)F=ma

here the net force is

F=F_{1}+F_{2}\\ \\ F=0.111\times 0.202+0.116\times 0.202\\ \\ F=0.0459N

in three significant figure